OpenStudy (gabylovesyou):

A firecracker shoots up from a hill 160 feet high with an initial speed of 105 feet per second. Using the formula H(t) = −16t2 + vt + s, determine how long it will take the firecracker to hit the ground. 7.5 seconds 7.8 seconds 8.1 seconds 8.3 seconds

OpenStudy (gabylovesyou):

@primeralph

OpenStudy (primeralph):

okay, do you know calculus?

OpenStudy (goformit100):

Madam is it a Physics question or Mathematics Question ?

OpenStudy (primeralph):

the whatever will be at the ground twice: when it leaves and when it falls. So, you will have two answers: the larger one is what you want. All you have to do is set the equation = 0, but for v, put160

OpenStudy (gabylovesyou):

calculus ? no im in algebra 1 @primeralph and math @goformit100

OpenStudy (primeralph):

let me see what you can come up with

OpenStudy (nincompoop):

you don't need calculus for this problem

OpenStudy (primeralph):

@nincompoop in real life, stuff don't work out so easily, except in algebra. That's why I asked.

OpenStudy (gabylovesyou):

im confused ... what do i do ?

OpenStudy (primeralph):

solve for t

OpenStudy (primeralph):

@Denebel she doesn't seem much like a physics person to me............

OpenStudy (nincompoop):

you're going to get a wrong answer if your v is set to 150

OpenStudy (gabylovesyou):

...... i get a decimal

OpenStudy (primeralph):

yeah, what's it?

OpenStudy (gabylovesyou):

10.342 and -0.96694

OpenStudy (gabylovesyou):

@primeralph

OpenStudy (nottim):

i think the reasonable one makes sense in that situation.

OpenStudy (anonymous):

The formula is H(t) = -16t^2 + 105t+160 = 0 (because v=105 as stated in your problem). You would also want to disregard any negative numbers because it is not possible to have a negative time

OpenStudy (primeralph):

errrr, but them answers ain't in the options

OpenStudy (anonymous):

If you solve h(t) = -16t^2 + 105t+160, it matches one of the choices (with rounding)

OpenStudy (primeralph):

@Denebel , not really.........

OpenStudy (gabylovesyou):

.-.

OpenStudy (nincompoop):

okay we're wasting a lot of time here...

OpenStudy (primeralph):

@Gabylovesyou you'll have to go with 10.34.... even though it isn't in the options

OpenStudy (nincompoop):

@Gabylovesyou summon me to help :) I don't want to be rude by taking over

OpenStudy (primeralph):

@poop, do your thing, time is running out.

OpenStudy (anonymous):

You need to use the quadratic formula \[\frac{ -b \pm \sqrt{b ^{2}+4ac} }{ 2a } = \frac{ -105\pm \sqrt{105^{2}-4(-16)(160} }{ 2(-16) }\] Plug this in your calculator. You will get a positive and negative answer. then disregard the negative one

OpenStudy (anonymous):

@goformit100: its a dynamics question :)

OpenStudy (nincompoop):

you know this requires several steps and one other solution is to use the quadratic formula set H(t) to 0 therefore you will have the formula 0 = -16t^2 + 105t+160 since you have a negative as your leading coefficient, you need to first multiply both sides by -1 giving you 0 = 16t^2 - 105t -160 then use the quadratic formula. this also can be solved by completing the square.

OpenStudy (primeralph):

oh, twas 105 not 150........ either way the answer ain't in the options

OpenStudy (nincompoop):

ya I told you from the start that you will get the wrong answer if you set your v to 150

OpenStudy (nincompoop):

-16 t^2+105 t+160 = 0 Multiply both sides by a term to make the coefficient of t^2 positive. Multiply both sides by -1: 16 t^2-105 t-160 = 0 Using the quadratic formula, solve for t. t = (105±sqrt((-105)^2-4×16 (-160)))/(2×16) = (105±sqrt(11025+10240))/32 = (105±sqrt(21265))/32: Answer: | t = (105+sqrt(21265))/32 or t = (105-sqrt(21265))/32

OpenStudy (anonymous):

As @Denebel said , Use quadratic formula ... Remember that the answer will be only da posivtive value

OpenStudy (anonymous):

Positive *

OpenStudy (nincompoop):

but since w're talking about time here and time doesn't go back, unless you want to be Einstein, then obtain the positive value.

OpenStudy (gabylovesyou):

whats the answer?

OpenStudy (nincompoop):

plug into your calculator

OpenStudy (nincompoop):

do you know how to complete the square yet, @Gabylovesyou

OpenStudy (primeralph):

−16t2 + 105t + 160 = 0 @Gabylovesyou please solve, as you solved earlier.

OpenStudy (nincompoop):

don't forget the caret sign, @primeralph ^ ^ ^ ^ ^ ^ ^ called caret

OpenStudy (gabylovesyou):

WUT? is it C?

OpenStudy (nincompoop):

the answer is in the option when it is rounded off to the first decimal

OpenStudy (gabylovesyou):

it has to be in the answers ... .-.

OpenStudy (gabylovesyou):

@nincompoop is it C?

OpenStudy (primeralph):

no, poop is right, it is among them.

OpenStudy (nincompoop):

look at this link and tell me if you see http://www.wolframalpha.com/input/?i=0+%3D+-16t%5E2+%2B+105t%2B160 the value very close to the one of the options

OpenStudy (primeralph):

@Gabylovesyou what type of calculator do you use?

OpenStudy (gabylovesyou):

is it B ?

OpenStudy (gabylovesyou):

@primeralph i use wolfram

OpenStudy (primeralph):

yeah, B. I mean handheld.

OpenStudy (gabylovesyou):

lol finally! ITS B!!!!!!!!!!!!!! right @nincompoop